OpenStudy (anonymous):

y = -0.57x^2 + 6.79x - 16.68 how to put in general form showing all steps?

OpenStudy (anonymous):

what is "general form"? looks like it is already in standard form

OpenStudy (anonymous):

idk... some people are saying its standard form and others are saying its general so im way confussed..

OpenStudy (anonymous):

wonder if this is the genral form f(x) = a(x - h)2 + k.......i think it's already in gen form....is it about conics?

OpenStudy (anonymous):

then i thought i was in general form already to until i saw ppl saying it was standard.. i think this one is standard f(x) = a(x - h)2 + k and this one is general form f(x) = ax2 + bx + c so it would be as my equation and i would have to change it to standard?

OpenStudy (anonymous):

may be...u kno how to solve it r8

OpenStudy (anonymous):

if it's standard

OpenStudy (anonymous):

no i dont ..

OpenStudy (anonymous):

um ok......if u have a.x2 + bx + c...... a(x^2 + kx)+c a(x^2+ kx+p - p)+c ..... a(x^2 +kx + p) +c-ap.... a(x - h)2 + k

OpenStudy (anonymous):

here is the answer i got ... y = -0.57 (x - 6.79/ 0.57) ^2 + (46.1041/ 2.28) - 16.68 i just need to know if its right or not?

OpenStudy (anonymous):

um....m lazy....

OpenStudy (anonymous):

haha thats fine! thanks anyway :)

OpenStudy (anonymous):

wait you want it in vertex form? \[y=a(x-h)^2+k\]?

OpenStudy (anonymous):

we can do that no problem, if that is what you want

OpenStudy (anonymous):

oh wait, looks like you have already done it, but there is a mistake

OpenStudy (anonymous):

well i can show you my work if that would help?

OpenStudy (anonymous):

\[y = -0.57x^2 + 6.79x - 16.68\] \[y=-.57(x^2-\frac{6.79}{.57}x)-16.68\] \[y=-.57(x-\frac{6.79}{1.14})^2-\text{something}\]

OpenStudy (anonymous):

you jumped a step. you have to factor out the -.57 first, then take half the coefficient of the middle term to complete the square

OpenStudy (anonymous):

these decimals are as annoying as can be, but i guess you are supposed to approximate. it doesn't really make sense to have one decimal divide by another, so you can either use a calculator and divide or else you could write \[y=-.57(x-\frac{679}{114})^2+\text{something}\] and you can find "something" by replacing \(x\) by \(\frac{679}{114}\) in the original expression to see what you get

OpenStudy (anonymous):

my guess is you are supposed to use a calculator to get it

OpenStudy (anonymous):

for the record the answer is \[y=-.57(x-\frac{679}{114})^2+\frac{80737}{22800}\]

OpenStudy (anonymous):

general idea is that if you start with \[y=ax^2+bx+c\] you go to \[y=a(x^2+\frac{b}{a}x)+c\] to \[y=a(x+\frac{b}{2a})^2+\text{something}\] and the something is what you get when you replace \(x\) by \(-\frac{b}{2a}\)

OpenStudy (anonymous):

yeah i know their annoying i had to do an activity to where i had to graph a parabola and this is the equation i had to work with sadly .. but thanks, you helped me alot!

OpenStudy (anonymous):

if you want to graph it, take a look here http://www.wolframalpha.com/input/?i=y+%3D+-0.57x^2+%2B+6.79x+-+16.68+ also if you want to see it in vertex form, look at the very bottom where it says what the max is and what value of x gives it to you match it up to what i wrote above, and you will see it

OpenStudy (anonymous):

oh okay well i did the graph